# On the equivalence of all notions of generalized derivations whose domain is a C-algebra

### v1

@@ Abstract.
    $F$ is a generalized derivation of the first type if there exists a derivation $d:\mathcal{A}\to\mathcal{M}$ satisfying $F(ab)=F(a)b+ad(b)$ for all $a,b\in\mathcal{A}$.
@@
    $F$ is a generalized derivation of the second type if there exists an element $\xi\in\mathcal{M}^{**}$ satisfying $F(ab)=F(a)b+aF(b)-a\xi b$ for all $a,b\in\mathcal{A}$.
@@
    $F$ is a generalized derivation of the third type if there exist two (non-necessarily linear) mappings $G,H:\mathcal{A}\to\mathcal{M}$ satisfying $F(ab)=G(a)b+aH(b)$ for all $a,b\in\mathcal{A}$.
There are examples showing that these three definitions are not, in general, equivalent. Despite that the first two notions are well studied when $\mathcal{A}$ is a C-algebra, it is not known if the three notions are equivalent under these special assumptions. In this note we prove that every generalized derivation of the third type whose domain is a C-algebra is automatically continuous. We also prove that every (continuous) generalized derivation of the third type from a C-algebra $\mathcal{A}$ into a general Banach $\mathcal{A}$-bimodule is a generalized derivation of the first and second type. In particular, the three notions coincide in this case. We also explore the possible notions of generalized Jordan derivations on a C-algebra and establish some continuity properties for them.
@@ 1. Introduction
Since early nineties, several notions of generalized derivations from an algebra to a bimodule have been considered in the literature, all of them built on an appropriate weak version of the proper definition of derivation. Derivations are among the most studied maps in the literature. Recall that a linear mappings $d$ from an associative algebra $\mathcal{A}$ to an $\mathcal{A}$-bimodule $\mathcal{M}$ is called a derivation (respectively, a Jordan derivation) if it satisfies Leibniz’ rule
@@
A simple example of derivation can be given by fixing an element $x_{0}\in\mathcal{M}$ and defining the mapping $d_{x_{0}}:\mathcal{A}\rightarrow\mathcal{M}$ given by the commutator associated with $x_{0}$, that is, $d_{x_{0}}(a)=[a,x_{0}]=ax_{0}-x_{0}a$. Such a derivation is called inner derivation.
The available literature contains at least three different uses of the term “generalized derivation”. The first one, in chronological order, appears in a paper by Brešar published in 1991 [4]. Keeping the notation above, a linear mapping $G:\mathcal{A}\to\mathcal{M}$ is a generalized derivation of the first type if there exists a derivation $d:\mathcal{A}\to\mathcal{M}$ such that the identity
@@
holds for all $a,b\in\mathcal{A}$. In the original definition, $G$ is only assumed to be additive. Note that (1.1) is equivalent to say that $G-d$ is a left multiplier, that is,
@@
In this case, the structure of a right $\mathcal{A}$-ideal on $\mathcal{M}$ is sufficient. Right multipliers can be similarly defined. This first notion is the one employed, for example, in the study by Heller, Miller, Pysiak and Sasin connecting differential geometry (connection, curvature, etc.), generalized derivations, and general relativity [18].
Keeping the chronological order, the second notion of generalized derivation was introduced by Nakajima in 1999 (cf. [36, $(1.3)$]). A reformulation of this notion was considered by Leger and Luks in [32]. This is the notion employed by Alaminos, Brešar, Extremera and Villena in the study of bounded linear operators preserving zero products (see [1, Definition 4.1]), and for example in [2, 11, 34, 35] and [40]. A generalized derivation of the second type from a Banach algebra $\mathcal{A}$ into a Banach $\mathcal{A}$-bimodule $\mathcal{M}$ is a linear mapping $G:\mathcal{A}\to\mathcal{M}$ for which there exists $\xi\in\mathcal{M}^{**}$ satisfying
@@
Every derivation is a generalized derivation, though the class of generalized derivations is strictly wider than the set of derivations (e.g. for each $a\in\mathcal{A}$, the mapping $x\mapsto x\circ a=\frac{1}{2}(ax+xa)$ is a generalized derivation which is not a derivation). Let us observe that if $\mathcal{A}$ and $\mathcal{M}$ are unital with unit $\mathbf{1}$, module products of the from $a\xi\mathbf{1}$ and $\mathbf{1}\xi b$ lie in $\mathcal{M}$ for all $a,b\in\mathcal{A}$, and hence the left and right multiplication operators $L_{\mathbf{1}\xi},R_{\xi\mathbf{1}}:a\mapsto\xi a,a\xi$ define two bounded linear operators from $\mathcal{A}$ to $\mathcal{M}$ (i.e., $\mathbf{1}\xi$ and $\xi\mathbf{1}$ behave like a multiplier). In this case, the mapping $d=G-L_{\mathbf{1}\xi}:\mathcal{A}\to\mathcal{M}$ is a derivation and $G(ab)=G(a)b+ad(b)$ for all $a,b\in\mathcal{A}$. Therefore, every generalized derivation of the second type is generalized derivation of the first type.
If $\mathcal{A}$ is a unital algebra, and $G:\mathcal{A}\to\mathcal{M}$ is a generalized derivation of the first type with associated derivation $d$. Since $G-d$ is a left multiplier, we have $(G-d)(a)=(G-d)(\mathbf{1})a$ for all $a\in\mathcal{A}$, and thus
@@
Furthermore, every bounded left multiplier $L$ from a general C-algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule $\mathcal{M}$, is of the form $L(a)=\xi a,$ where $\xi\in M^{**}$ satisfies $\xi\mathcal{A}\subseteq\mathcal{A}$ (cf. [1, 2]). As before, we see that in this case, continuous generalized derivations of the first and second types from $\mathcal{A}$ to $\mathcal{M}$ coincide, thanks to the existence of bounded approximate units. It is important to note that continutiy has been assumed to establish the equivalence between generalized derivations of the first two types from a general C-algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule. It should be also mentioned that every generalized derivation of the second type from a C-algebra into a Banach bimodule is automatically continuous (cf. [25, Proposition 2.1]). Hence, when the domain is a C-algebra $\mathcal{A}$, the first two types of (continuous) generalized derivations agree.
To introduce the third notion of generalized derivation, it is necessary to revisit the definition of a ternary derivation or a $3$-tuple behaving like a derivation, introduced by Jimenéz-Gestal and Pérez-Izquierdo in [26] and Shestakov in [40, 41]. However, here we relax the linearity assumptions on the last two maps in the 3-tuple.
@@
Let $\mathcal{A}$ be an algebra, and let $\mathcal{M}$ be an $\mathcal{A}$-bimodule. A ternary derivation from $\mathcal{A}$ to $\mathcal{M}$ is a 3-tuple $(F,G,H)$, where $G,H:\mathcal{A}\rightarrow\mathcal{M}$ are two (non-necessarily linear) mappings, and $F:\mathcal{A}\rightarrow\mathcal{M}$ is a linear map satisfying $F(ab)=G(a)b+aH(b)$ for all $a,b\in\mathcal{A}$. We shall also say that $F:\mathcal{A}\rightarrow\mathcal{M}$ is a ternary derivation with associated mappings $G,H:\mathcal{A}\rightarrow\mathcal{M}$. Since the term “ternary derivation” is also employed in another settings (like in the case of JB-triples) with another meaning, we shall better say that the triplet $(F,G,H)$ behaves like a derivation. In this case the mapping $F$ is called a generalized derivation of the third type.
@@
Komatsu and Nakajima presented in [31] a detailed study on the relations among the notions of generalized derivations of first, second and third type and their other formal properties, mainly in the unital case and from an algebraic perspective. It is perhaps worth to recall some basic connections. Suppose $F:\mathcal{A}\to\mathcal{M}$ is a generalized derivation of the second type satisfying (1.2). The expression
$F(ab)=(F-R_{\xi})(a)b+aF(b)=F(a)b+a(F-L_{\xi})(b),$
is valid for all $a,b\in\mathcal{A}$, and hence $(F,F-R_{\xi},F)$ and $(F,F,F-L_{\xi})$ behave like derivations. That is, every generalized derivation of the second type is a generalized derivation of the third type. Conversely, it was already observed by Shestakov in [40, Lemma 1] that if $\mathcal{A}$ is a unital associate algebra, every generalized derivation of the third type $F:\mathcal{A}\to\mathcal{A}$ is generalized derivation of the second type (see also [31, Lemma 4.1 and Corollary 4.5]). The argument works in our general setting. Suppose $(F,G,H)$ is a $3$-tuple of mappings from a unital associative algebra $\mathcal{A}$ into a unital $\mathcal{A}$-bimodule behaving like a derivation, with $F$ being linear. As we shall see in Lemma 2.2, both maps $G$ and $H$ are linear and satisfy $G(a)=F(a)-aH(\mathbf{1})$ and $H(a)=F(a)-G(\mathbf{1})a$ for all $a\in\mathcal{A}$. Therefore it follows that
@@
However, if we relax the assumptions that $\mathcal{A}$ and the $\mathcal{A}$-bimodule are both unital, it is not clear whether every generalized derivation of the third type is of the second type. This naturally gives rise to a question: are these three notions of generalized derivation defined from a nonunital algebra $\mathcal{A}$ into a $\mathcal{A}$-bimodule equivalent with each other? In this paper, we shall give an affirmative answer to this question when $\mathcal{A}$ is a general C-algebra. In fact, this problem is equivalent to a question on automatic continuity.
@@
([1, Theorem 4.5], [7, Proposition 4.3], [2, Theorem 2.11]) Let $T:\mathcal{A}\to X$ be a bounded linear operator from a C-algebra to an essential Banach $\mathcal{A}$-bimodule. Then the following statements are equivalent:
@@ 3. $(c)$
In case of a linear mapping acting on a von Neumann algebra $\mathcal{W}$ and behaving like a generalized derivations at certain points of the domain, continuity becomes an inherent property. More concretely, let $T:\mathcal{W}\to\mathcal{W}$ be a linear mapping on a von Neumann algebra. Suppose that for each $a,b,c$ in any commutative von Neumann subalgebra $\mathcal{B}\subseteq\mathcal{W}$ with $ab=bc=0$ we have $aT(b)c=0$. Then $T$ is automatically continuous [11, Theorem 2.12 and Corollary 2.15]. Furthermore, the above statements $(a)$$(c)$ are also equivalent to the next:
@@ 2. $(e)$
The list of studies exploring the automatic continuity of derivations and related operators is quite wide. The pioneering theorems by Sakai [43] and Ringrose [38] prove that every derivation on a C-algebra or from a C-algebra $\mathcal{A}$ into a Banach $\mathcal{A}$-bimodule is automatically continuous, respectively. Hou and Ming [24] proved that if $\mathcal{X}$ is a simple Banach space, and $\sigma,\tau:B(\mathcal{X})\rightarrow B(\mathcal{X})$ are surjective and continuous at 00, then every $(\sigma,\tau)$-derivation from $B(\mathcal{X})$ into itself is continuous. Recall that, if $\mathcal{A}$ in an algebra and $\sigma,\tau:\mathcal{A}\rightarrow\mathcal{A}$ are two mappings, a $(\sigma,\tau)$-derivation on $\mathcal{A}$ is a linear mapping $d:\mathcal{A}\rightarrow\mathcal{A}$ satisfying
@@
Let us explain how, in the setting of C-algebras and essential Banach bimodules, the problem of determining whether every generalized derivation of the third type is of the second type is a problem of automatic continuity. Suppose $F,G,H$ are three mappings from a C-algebra $\mathcal{A}$ to an essential $\mathcal{A}$-bimodule $\mathcal{M}$, such that the triplet $(F,G,H)$ behaves like a derivation, and let us assume that $F$ is continuous. For arbitrary $a,b,c\in\mathcal{A}_{sa}$ with $ab=bc=0$. Choose, via functional calculus, a decomposition of $b$ in the form $b=b^{+}-b^{-}$, with $b^{+},b^{-}\geq 0$, $b^{+}b^{-}=0,$ $ab^{+}=ab^{-}=0$, $b^{+}c=0$, and $b^{-}c=0.$ By taking $d=\left(b^{+}\right)^{\frac{1}{2}}+i\left(b^{-}\right)^{\frac{1}{2}}$, we have $ad=dc=0$, $d^{2}=b$ and
@@
Theorem 1.2 implies that $F$ is a generalized derivation of the second type. So it suffices to prove that every generalized of the third type from a C-algebra into an essential Banach bimodule is continuous automatically. In Proposition 2.5 we prove that every continuous generalized derivation of the third type from a C-algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule is a generalized derivation of the first and second type, that is, we do not need to assume that the bimodule is essential. Theorem 2.8 completes the picture by showing that every generalized derivation of the third type from a C-algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule $\mathcal{M}$ is continuous. Consequently, every generalized derivation of the third type from $\mathcal{A}$ to $\mathcal{M}$ is a generalized derivation of the second type (and of course, of the first type).
Generalized Jordan derivations of the second type from a C-algebra $\mathcal{A}$ into a Banach $\mathcal{A}$-bimodule $\mathcal{M}$ have been already considered, for example, in [2, 6, 7]. A linear mapping $F:\mathcal{A}\rightarrow\mathcal{M}$ is a generalized Jordan derivation of the second type if there exists an element $\xi\in\mathcal{M}^{**}$ satisfying
@@
Theorem 3.1 shows that if $\mathcal{A}$ is a C-algebra, every generalized Jordan derivation of the third type $F$ from $\mathcal{A}$ into a Banach $\mathcal{A}$-bimodule $\mathcal{M}$ is continuous. Moreover, if $\mathcal{M}$ is essential, $F$ is a generalized derivation of the second type; and, in certain cases (for example, when $\mathcal{M}=\mathcal{A}^{*}$ or $\mathcal{M}=\mathcal{A}^{*}$), generalized Jordan derivations of the third type will be generalized Jordan derivation of the second type whose associated element in $\mathcal{M}^{**}$ will commute with all elements in $\mathcal{A}$ (see Proposition 3.2).

### v2

@@ Abstract.
    $F$ is a generalized derivation of the first type if there exists a derivation $d:\mathcal{A}\to\mathcal{M}^{**}$ satisfying $F(ab)=F(a)b+ad(b),$ for all $a,b\in\mathcal{A}$.
@@
    $F$ is a generalized derivation of the second type if there exists an element $\xi\in\mathcal{M}^{**}$ satisfying $F(ab)=F(a)b+aF(b)-a\xi b,$ for all $a,b\in\mathcal{A}$.
@@
    $F$ is a generalized derivation of the third type if there exist two (non-necessarily linear) mappings $G,H:\mathcal{A}\to\mathcal{M}$ satisfying $F(ab)=G(a)b+aH(b),$ for all $a,b\in\mathcal{A}$.
These three types of maps are not, in general, equivalent. Although the first two notions are well studied when $\mathcal{A}$ is a C-algebra, their connections with the third one have not yet been explored. In this note we prove that every generalized derivation of the third type from a C-algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule $\mathcal{M}$ is automatically continuous. We also show that every (continuous) generalized derivation of the third type from $\mathcal{A}$ to $\mathcal{M}$ is a generalized derivation of the first and second type. Consequently, the three notions coincide in this case. We also explore some concepts of generalized Jordan derivations on a C-algebra and establish some continuity properties for them.
@@ 1. Introduction
Several notions of generalized derivations from an algebra $\mathcal{A}$ to a bimodule $\mathcal{M}$ have been considered in the literature since early nineties. Derivations are among the most studied maps in the literature. Recall that a linear mapping $d$ from an associative algebra $\mathcal{A}$ to an $\mathcal{A}$-bimodule $\mathcal{M}$ is called a derivation (respectively, a Jordan derivation) if it satisfies Leibniz’ rule, that is,
@@
For example, if we fix an element $x_{0}\in\mathcal{M},$ the mapping $d_{x_{0}}:\mathcal{A}\rightarrow\mathcal{M}$ given by $d_{x_{0}}(a)=[a,x_{0}]=ax_{0}-x_{0}a$ is a derivation. Derivations which are expressed as finite sums of maps of the form $d_{x_{0}}$ are called inner derivations.
The available literature contains at least three different uses of the term “generalized derivation”. The first one, in chronological order, appears in a paper by Brešar published in 1991 (see [4]). Keeping the notation above, a linear mapping $G:\mathcal{A}\to\mathcal{M}$ is a generalized derivation of the first type if there exists a derivation $d:\mathcal{A}\to\mathcal{M}$ such that the identity
@@
holds for all $a,b\in\mathcal{A}$. In the original definition, $G$ is only assumed to be additive. Along this paper, we shall say that a linear mapping $G$ from a Banach algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule $\mathcal{M}$ is a generalized derivation of the first type if there exists a derivation $d:\mathcal{A}\to\mathcal{M}^{**}$ such that the identity (1.1) holds for all $a,b\in\mathcal{A}$. Note that the identity in (1.1) is equivalent to say that $G-d:\mathcal{A}\to\mathcal{M}^{**}$ is a left multiplier, that is,
@@
Right multipliers can be similarly defined. This first notion is the one employed, for example, in the study by Heller, Miller, Pysiak, and Sasin connecting differential geometry, generalized derivations, and general relativity [19].
Keeping the chronological order, the second definition of a generalized derivation was introduced by Nakajima in 1999 (cf. [37, $(1.3)$]). A reformulation of this notion was considered by Leger and Luks in [33]. This is the notion employed by Alaminos, Brešar, Extremera, and Villena in the study of bounded linear operators preserving zero products (see [1, Definition 4.1]), and for example in [2, 11, 35, 36] and [41]. A generalized derivation of the second type from a Banach algebra $\mathcal{A}$ into a Banach $\mathcal{A}$-bimodule $\mathcal{M}$ is a linear mapping $G:\mathcal{A}\to\mathcal{M}$ for which there exists $\xi\in\mathcal{M}^{**}$ satisfying
@@
Every derivation is a generalized derivation of the second type, though the class determined by the latter maps is strictly wider (e.g. for each non-zero $a\in\mathcal{A}$, the mapping $x\mapsto x\circ a=\frac{1}{2}(ax+xa)$ is a generalized derivation of the second type on $\mathcal{A}$ which is not a derivation). Let us observe that if $\mathcal{A}$ is unital, with unit $\mathbf{1}$, module products of the from $a\xi\mathbf{1}$ and $\mathbf{1}\xi b$ lie in $\mathcal{M}$ for all $a,b\in\mathcal{A}$, and hence the left and right multiplication operators $L_{\mathbf{1}\xi},R_{\xi\mathbf{1}}:a\mapsto\mathbf{1}\xi a,a\xi\mathbf{1}$ define two bounded linear operators from $\mathcal{A}$ to $\mathcal{M}$ (i.e., $\mathbf{1}\xi$ and $\xi\mathbf{1}$ behave like a multiplier). In this case, the mapping $d=G-L_{\mathbf{1}\xi}:\mathcal{A}\to\mathcal{M}$ is a derivation and $G(ab)=G(a)b+ad(b)$ for all $a,b\in\mathcal{A}$. Therefore, every generalized derivation of the second type is a generalized derivation of the first type.
Suppose now that $\mathcal{A}$ is a unital algebra, and $G:\mathcal{A}\to\mathcal{M}$ is a generalized derivation of the first type with associated derivation $d$. Since $G-d$ is a left multiplier, we have $(G-d)(a)=(G-d)(\mathbf{1})a$ for all $a\in\mathcal{A}$, and thus
@@
Furthermore, every bounded left multiplier $L$ from a C-algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule $\mathcal{M}$, is of the form $L(a)=\xi a,$ where $\xi\in M^{**}$ satisfies $\xi\mathcal{A}\subseteq\mathcal{A}$ (cf. [1, 2]). Therefore, every continuous generalized derivation of the first type from $\mathcal{A}$ to $\mathcal{M}$ is a generalized derivation of the second type. Conversely, if $G:\mathcal{A}\to\mathcal{M}$ is a generalized derivation of the second type with respect to an element $\xi\in\mathcal{M}^{**}$, the mapping $d=G-L_{\xi}:\mathcal{A}\to\mathcal{M}^{**}$ is a derivation, and $G(ab)=G(a)b+ad(b)$ for all $a,b\in\mathcal{A}$. We have thus shown that continuous generalized derivations of the first and second type from $\mathcal{A}$ to $\mathcal{M}$ coincide. It is important to note that continuity has been assumed to establish the equivalence between generalized derivations of the first two types from a C-algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule. It should also be mentioned that every generalized derivation of the second type from a C-algebra into a Banach bimodule is automatically continuous (cf. [26, Proposition 2.1]).
The third notion of generalized derivation is motivated by a definition introduced by Jimenéz-Gestal and Pérez-Izquierdo in [27], and employed by Shestakov in [41, 42]. However, here we relax some linearity assumptions included in the original statement.
@@
Let $\mathcal{A}$ be an algebra, and let $\mathcal{M}$ be an $\mathcal{A}$-bimodule. A ternary derivation from $\mathcal{A}$ to $\mathcal{M}$ is a $3$-tuple $(F,G,H)$, where $G,H:\mathcal{A}\rightarrow\mathcal{M}$ are two (non-necessarily linear) mappings, and $F:\mathcal{A}\rightarrow\mathcal{M}$ is a linear map satisfying $F(ab)=G(a)b+aH(b)$ for all $a,b\in\mathcal{A}$. We shall also say that $F:\mathcal{A}\rightarrow\mathcal{M}$ is a ternary derivation with associated mappings $G,H:\mathcal{A}\rightarrow\mathcal{M}$. Since the term “ternary derivation” is also employed in other settings with another meaning (like in the case of JB-triples), we shall better say that the triplet $(F,G,H)$ behaves like a derivation. In this case, the mapping $F$ is called a generalized derivation of the third type.
@@
Komatsu and Nakajima presented in [32] a detailed study on the relationships among the different notions of generalized derivations of first, second, and third type, and their formal properties; but mainly in the unital case and from an algebraic perspective. It is perhaps worth recalling some basic connections. Suppose $F:\mathcal{A}\to\mathcal{M}$ is a generalized derivation of the second type satisfying (1.2). By assuming that $\mathcal{A}$ is unital, it follows that the element $\mathbf{1}\xi\mathbf{1}=\mathbf{1}G(\mathbf{1})\mathbf{1}\in\mathcal{M}$. The expression
$F(ab)=(F-R_{\mathbf{1}\xi\mathbf{1}})(a)b+aF(b)=F(a)b+a(F-L_{\mathbf{1}\xi\mathbf{1}})(b),$
is valid for all $a,b\in\mathcal{A}$, and hence the triplets $(F,F-R_{\mathbf{1}\xi\mathbf{1}},F)$ and $(F,F,F-L_{\mathbf{1}\xi\mathbf{1}})$ behave like derivations. That is, every generalized derivation of the second type is a generalized derivation of the third type. Conversely, it was already observed by Shestakov in [41, Lemma 1] that if $\mathcal{A}$ is a unital associate algebra, every generalized derivation of the third type $F:\mathcal{A}\to\mathcal{A}$ is a generalized derivation of the second type (see also [32, Lemma 4.1 and Corollary 4.5]). The argument works in our general setting too. Suppose $(F,G,H)$ is a $3$-tuple of mappings from a unital associative algebra $\mathcal{A}$ into a unital $\mathcal{A}$-bimodule, $\mathcal{M},$ behaving like a derivation, with $F$ being linear. As we shall see in Lemma 2.2, the mappings $G\mathbf{1}$ and $\mathbf{1}H$ are linear and satisfy $G(a)\mathbf{1}=F(a)-a\mathbf{1}H(\mathbf{1})$ and $\mathbf{1}H(a)=F(a)-G(\mathbf{1})a,$ for all $a\in\mathcal{A}$. Therefore, it follows that
@@
However, if we relax the assumption that $\mathcal{A}$ is unital, it is not clear whether every generalized derivation of the third type is of the second type. This naturally gives rise to a question: are these three notions of generalized derivations from a (non-necessarily unital) C-algebra $\mathcal{A}$ into an $\mathcal{A}$-bimodule equivalent with each other? In this paper, we shall give an affirmative answer to this question. In fact, this problem is equivalent to a question on automatic continuity.
@@
([1, Theorem 4.5], [7, Proposition 4.3], [2, Theorem 2.11]) Let $T:\mathcal{A}\to\mathcal{M}$ be a bounded linear operator from a C-algebra to an essential Banach $\mathcal{A}$-bimodule. Then the following statements are equivalent:
@@ 3. $(c)$
Recall that a Banach $\mathcal{A}$-bimodule $\mathcal{M}$ is called essential if the linear span of the set $\{axb:a,b\in A,x\in\mathcal{M}\}$ is dense in $\mathcal{M}$. It is worth noting that in case that $\mathcal{A}$ is a Banach algebra admitting a bounded approximate unit $(u_{j})_{j}$ and $\mathcal{M}$ is essential, we can deduce from the celebrated Cohen factorization theorem (see [20, Theorem 32.22] and [8, Theorem in page 108, § 1.15]) that, for each $x\in\mathcal{M}$ the nets $(u_{j}x)_{j}$ and $(xu_{j})_{j}$ converge in norm to $x$.
In case of a linear mapping acting on a von Neumann algebra $\mathcal{W}$ and behaving like a generalized derivation at certain points of the domain, continuity becomes an automatic property. More concretely, let $T:\mathcal{W}\to\mathcal{W}$ be a linear mapping on a von Neumann algebra. Suppose that for each $a,b,c$ in any commutative von Neumann subalgebra $\mathcal{B}\subseteq\mathcal{W}$ with $ab=bc=0$ we have $aT(b)c=0$. Then $T$ is automatically continuous [11, Theorem 2.12 and Corollary 2.15]. Furthermore, the above statements $(a)$$(c)$ are also equivalent to the next:
@@ 2. $(e)$
The list of studies exploring the automatic continuity of derivations and related operators is quite wide. The pioneering theorems by Sakai [44] and Ringrose [39] prove that every derivation on a C-algebra and every derivation from a C-algebra $\mathcal{A}$ into a Banach $\mathcal{A}$-bimodule is automatically continuous. Hou and Ming [25] proved that if $\mathcal{X}$ is a simple Banach space, and $\sigma,\tau:B(\mathcal{X})\rightarrow B(\mathcal{X})$ are surjective and continuous at 00, then every $(\sigma,\tau)$-derivation from $B(\mathcal{X})$ into itself is continuous. Recall that, if $\mathcal{A}$ is an algebra and $\sigma,\tau:\mathcal{A}\rightarrow\mathcal{A}$ are two mappings, a $(\sigma,\tau)$-derivation on $\mathcal{A}$ is a linear mapping $d:\mathcal{A}\rightarrow\mathcal{A}$ satisfying
@@
Let us explain how, in the setting of C-algebras and essential Banach bimodules, the problem of determining whether every generalized derivation of the third type is of the second type is a problem of automatic continuity. Suppose $F,G,H$ are three mappings from a C-algebra $\mathcal{A}$ to an essential $\mathcal{A}$-bimodule $\mathcal{M}$, such that the triplet $(F,G,H)$ behaves like a derivation, and let us assume that $F$ is continuous. For arbitrary $a,b,c\in\mathcal{A}_{sa}$ with $ab=bc=0$, let us choose, via functional calculus, a decomposition of $b$ in the form $b=b^{+}-b^{-}$, with $b^{+},b^{-}\geq 0$, $b^{+}b^{-}=0,$ $ab^{+}=ab^{-}=0$, $b^{+}c=0$, and $b^{-}c=0.$ By taking $d=\left(b^{+}\right)^{\frac{1}{2}}+i\left(b^{-}\right)^{\frac{1}{2}}$, we have $ad=dc=0$, $d^{2}=b,$ and
@@
Thus, Theorem 1.2 implies that $F$ is a generalized derivation of the second type. So it suffices to prove that every generalized derivation of the third type from a C-algebra into an essential Banach bimodule is continuous. In Proposition 2.5 we prove that every continuous generalized derivation of the third type from a C-algebra $\mathcal{A}$ to a (non-necessarily essential) Banach $\mathcal{A}$-bimodule is a generalized derivation of the first and second type. Theorem 2.8 completes the picture by showing that every generalized derivation of the third type from a C-algebra $\mathcal{A}$ to a Banach $\mathcal{A}$-bimodule $\mathcal{M}$ is continuous. Consequently, every generalized derivation of the third type from $\mathcal{A}$ to $\mathcal{M}$ is a generalized derivation of the second type (and of course, of the first type).
Generalized Jordan derivations of the second type from a C-algebra $\mathcal{A}$ into a Banach $\mathcal{A}$-bimodule $\mathcal{M}$ have already been considered (see, for example, [2, 6, 7]). A linear mapping $F:\mathcal{A}\rightarrow\mathcal{M}$ is a generalized Jordan derivation of the second type if there exists an element $\xi\in\mathcal{M}^{**}$ satisfying
@@
Theorem 3.1 below shows that if $\mathcal{A}$ is a C-algebra, every generalized Jordan derivation of the third type from $\mathcal{A}$ into a Banach $\mathcal{A}$-bimodule $\mathcal{M}$ is continuous. Moreover, if we additionally assume that $\mathcal{M}$ is essential, then $F$ is a generalized derivation of the second type; and in certain cases (for example, when $\mathcal{M}=\mathcal{A}$ or $\mathcal{M}=\mathcal{A}^{*}$) generalized Jordan derivations of the third type will be generalized Jordan derivations of the second type whose associated element in $\mathcal{M}^{**}$ will commute with all elements in $\mathcal{A}$ (see Proposition 3.2).